# Beyond Calibration: Assessing the Probabilistic Fit of Neural Regressors via Conditional Congruence

### v2

@@ 4.1 CCE with a Known Data-Generating Process (Synthetic Data)
ECE ($\downarrow$) NLL ($\downarrow$) $\overline{\mbox{CCE}}$ ($\downarrow$)
Gaussian NN 0.235 1.563 0.086 (0.009)
Stirn et al. [12] 0.148 1.404 0.100 (0.007)
Immer et al. [13] 0.165 1.207 0.041 (0.002)
Seitzer et al. [41] 0.194 1.389 0.068 (0.004)
4.2 CCE with an Unknown Data-Generating Process (Image Regression)
In the following section we demonstrate that CCE can effectively scale beyond low-dimensional data. To do so, we perform a series of image regression tasks. In real-world problems, the data generating process is not known. However, we show that CCE is effective at identifying increasing divergence between a model’s predictive distribution and test data when we apply growing levels of perturbations. We then provide a comparison of CCE, ECE, and NLL across a variety of datasets.
4.2.1 CCE Under Known Distribution Shift
We perform an image regression task where the goal is to accurately predict the number of people in an image using the ‘person’ class of the MS-COCO dataset [44]. The bounding boxes for each image have been translated into a single, discrete count [16]. This task allows us to train regressors with both continuous and discrete predictive distributions.
We train three different regression models, each with a different predictive distribution: a Poisson NN [14], $\beta$-Gaussian ($\beta=0.5$) [41] and DDPN [16]. Each network is trained with the same backbone, the pooled output from a ViT [45] (initialized from the vit-base-patch16-224-in21k checkpoint). Both $\beta$-Gaussian and DDPN are modern, highly flexible regressors, while Poisson is subject to equi-dispersion and has a more rigid parametric form. This diversity in models allows us to observe how well CCE captures increasing levels of data shift under a variety of predictive distributions.
For each model, we perform three types of data shift. First, we apply varying levels of Gaussian blur to the input image, which perturbs the marginal of the input, $P_{X}$, away from the training data. Second, we apply Gaussian noise to the marginal of the output, $P_{Y}$. Third, we apply the Mixup transform [42] to both the images and the labels, which is a perturbation to the conditional, $P_{Y|X}$. For each model, we perform five trials of the data perturbation strategy and evaluation, and plot the mean and standard deviation of the resulting $\overline{\mbox{CCE}}$ statistic. For each instance in the test data, we draw a single Monte Carlo sample from the predictive distribution ($\ell=1$).
Results are shown in Figure 3. We see that under all three data shift regimes, $\overline{\mbox{CCE}}$ smoothly increases as as perturbations of progressively larger magnitude are applied. Additionally, we observe that because CCE is a measure of conditional congruence, it can identify misalignment in both the conditional, $P_{Y|X}$, and marginal distributions, $P_{X}$ and $P_{Y}$. Additionally, CCE can even detect divergence when a model’s predictive distribution is less expressive, such as the Poisson network, which only has one parameter to represent predictive mean and variance.
4.2.2 Comparing CCE to ECE and NLL Across Datasets
Next we diagnose the behavior of CCE across a variety of tasks and datasets. In these experiments, there is no model that is a priori known to be best-fit, and the data generating process is unknown. Our goal is to compare and contrast the measures of probabilistic alignment provided by $\overline{\mbox{CCE}}$, ECE, and NLL.
The regression tasks studied include the COCO-People task introduced above, predicting a person’s age from an image (FG-Net [46], AAF [47]), and regressing a score describing how appealing an image is (EVA [48]). On all counting tasks (COCO-People, FG-Net, and AAF), we train a Poisson NN [14], a Negative Binomial NN [15], a DDPN [16], a vanilla Gaussian NN  [^2] Although modeling discrete data with a Gaussian distribution is not entirely proper, it is a common practice due to the distribution’s flexibility. [11], and modern Gaussian regressors such the $\beta$-modified Gaussian ($\beta=0.5$) [41], the naturally parameterized Gaussian [13], and the “faithful” heteroscedastic regressor [12]. For the EVA dataset, we only benchmark Gaussian models since the labels are continuous. All architecture, hardware, and hyperparameter details for these experiments can be found in Appendix A.2.
Tables 1 and 2 contain the $\overline{\mbox{CCE}}$, ECE and NLL for different models trained on COCO-People, FG-Net, AAF, and EVA. We compute $\overline{\mbox{CCE}}$ 5 times and report the mean and standard deviation of this statistic (in parentheses) across trials. CCE values were computed according to the description in Section 3.2.
In general, we observe broad agreement among $\overline{\mbox{CCE}}$, ECE, and NLL. The best model with respect to conditional congruence ($\overline{\mbox{CCE}}$) also exhibits the best (or near-best) marginal calibration (ECE) and average NLL. For example, on FG-Net, the Negative Binomial NN achieves the lowest error by all three measures. On AAF, Negative Binomial is once again the best performer on ECE and $\overline{\mbox{CCE}}$. Meanwhile, [13] scores best on NLL and $\overline{\mbox{CCE}}$ while coming in second with its ECE. These results are to be expected — since point-wise congruence implies marginal calibration, our metrics should tend toward agreement in the aggregate.
4.3 Diagnosing Point-wise Model Failures with CCE
While $\overline{\mbox{CCE}}$ is an effective overall descriptor of probabilistic fit, one of the greatest virtues of CCE is as a point-wise reliability measure. Specifically, by viewing the inputs that incur both the highest and lowest CCE values, we can gain insight into a model’s performance on specific regions of $\mathcal{X}$. Most importantly, this analysis can be done without access to labels. To do this, we take $\mathcal{S}$ to be the validation split of a dataset, and evaluate $\text{CCE}(f,\mathcal{S},\varphi,\cdot)$ at all inputs from the test split. The resultant values provide insight into the types of examples the model is best- and worst-fit on. We demonstrate this diagnostic process on two model-dataset pairs: [41]’s $\beta$-Gaussian that has been fit on the AAF dataset (see Figure 4) and the Poisson NN trained on COCO-People (see Figure 5).
In the first row of Figure 4, we plot the images from AAF with the lowest CCE values in our test data. These are images where the model’s learned conditional distributions are most congruent with the true data-generating distribution. In the second row, we plot the images with the highest CCE values, where the model’s distributions are least congruent. According to our measurement, the model struggles most in terms of probabilistic alignment when predicting the ages of young-adult women, while it appears to be more reliable when given images of middle-aged males, shedding light on possible biases learned by our model. While the AAF dataset is made up of around 55% female faces, each model’s feature extractor was pre-trained on much larger datasets where racial or gender bias likely exists [49].
Figure 5 reveals where the Poisson NN trained on COCO-People performs best (top row) and worst (bottom row) with respect to CCE. Interestingly, the Poisson model struggles to form distributions over severely high counts in crowded scenes. This may be due to the rigid equivariance assumption inherent to a Poisson model. If the true data-generating distribution in the region of “large crowds” is under- or over-dispersed, the Poisson NN would not be able to represent this fact. Thus, the model outputs distributions over $\mathcal{Y}$ that are incongruent with the ground truth. Overall, this analysis provides an effective framework for discovering point-wise incongruence for test examples.
0.0385
0.0388
0.0396
0.0402
0.0407
0.0897
0.0901
0.0903
0.0905
0.0918
Figure 4: Test images that incurred the lowest (first row) and highest (second row) CCE values for a $\beta$-Gaussian model [41] trained on the AAF dataset. CCE values indicated above each image. High values correspond to the largest distance between learned and actual conditional distributions.
0.0441
0.0441
0.0447
0.0453
0.0454
0.0758
0.0768
0.0772
0.0776
0.0800
Figure 5: Lowest and highest CCE values for a Poisson NN trained on COCO-People.
5 Conclusion
In this paper, we identify issues with current approaches used to evaluate the quality of predictive distributions produced by neural networks. We show that existing measurements that target calibration are inherently marginal and ignore conditional dependencies in the data. To address this, we introduce a stronger definition of probabilistic fit, called congruence, and propose Conditional Congruence Error (CCE) as a novel measure. CCE is a conditional, rather than marginal, statistic and can be computed point-wise for unseen test instances. We show CCE can effectively recover divergence between predictive distributions and test data and can be used to interpret failure cases of a model. Future work should study open questions about using CCE in the classification setting, using CCE as a regularizer during training, or incorporating point-wise CCE values into an active learning framework.
Appendix A SUPPLEMENTARY EXPERIMENTAL DETAILS
A.1 Derivation of the Marginal Distribution of $Y$ in Figure 1
@@
[…]
@@
A.2 Detailed Description of Experiments
In all experiments, the model that achieved the best loss on the validation set during training was selected for evaluation. This helped mitigate the overfitting that we observed with respect to NLL as training continued.
Data was split into train/val/test using a fixed random seed, allocating 70% to train, 10% to val, and 20% to test.
The AdamW optimizer [50] was used for all training jobs, with an initial learning rate of $10^{-3}$ that decayed to 0 on a cosine annealing schedule [51]. We used a weight decay value of $10^{-5}$. For image regression tasks, we applied the standard ImageNet normalization constants and augmented with the AutoAugment [52] transformation during training. Though we specify the feature encoder used for each dataset, we refer the reader to our source code  [^3] https://anonymous.4open.science/r/probcal-8294 as well as the original literature on each model to find exact details about the regression heads fit on top of each encoder.
The following subsections contain further details for the models we trained, organized by dataset. We also provide relevant licensing information and URLs for each dataset where applicable.
A.2.1 Synthetic Dataset
Following [16], the MLPs we trained for the synthetic dataset in 4.1 had layers of widths [128, 128, 128, 64]. Training ran for 200 epochs on a 2021 MacBook Pro CPU, with a batch size of 32.
We generated this dataset locally using the process defined by [16].
A.2.2 FG-Net, AAF, and EVA
@@
A.2.3 COCO-People
On COCO-People, we use the pooled output from a ViT backbone (initialized from the vit-base-patch16-224-in21k checkpoint [45, 54]), followed by a [384, 256] MLP with ReLU activations [55] for our feature encoder. This matches the setup used by [16]. Models were trained on an internal cluster of 2 NVIDIA GeForce RTX 4090 GPUs with an effective batch size of 256. Training ran for 30 epochs with BFloat16 Mixed Precision. All CCE evaluation was done on a single GPU of this cluster, which has 24GB of VRAM.
The COCO dataset from which [16] defines COCO-People is distributed with a CCBY 4.0 license. See https://cocodataset.org/#home.
Appendix B COMPUTING CCE — PRACTICAL CONSIDERATIONS
B.1 Algorithm to Compute CCE for a Model
Assuming an existing implementation of the MCMD (which can be easily derived from Equation 2 after a choice of kernel), we employ Algorithm 1 to compute CCE values for a given model $f$, embedder $\varphi$, and set of g.t. samples $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}$.
@@ B.1 Algorithm to Compute CCE for a Model
[…]
@@
for x in $\{x_{i}\}_{i=1}^{n}$ do
[…]
@@
For an implementation of this algorithm in Python, we refer the reader to our source code.
B.2 Accelerating the CCE Computation
Though we use loops for clarity in specifying Algorithm B.1, in practice, we vectorize all computations (including measuring the MCMD across a grid of values) to decrease runtime. In addition to this vectorization, we observe significant speed-ups (on the order of 15-20$\times$) when moving all matrix operations to a GPU.
The CCE computation can be accelerated even further (and made more numerically stable) by taking advantage of the structure of the Gram matrices involved in the MCMD (see Equation 2). Since $(\mathbf{K}_{X}+n\lambda\mathbf{I}_{n})$ and $(\mathbf{K}_{Y}+m\lambda^{\prime}\mathbf{I}_{m})$ are positive definite and symmetric, we compute their inverses via the Cholesky decomposition (which has half the cost of the typical LU decomposition performed by linear algebra libraries). Our provided code elegantly handles these details, yielding a performant and scalable CCE implementation.
B.3 Addressing Memory Limitations
As is common with kernel methods, the CCE (which relies on computing the MCMD for a collection of encoded inputs) grows rapidly in terms of memory requirements with the size of the data, exhibiting quadratic spatial complexity. In this work, we found that our hardware was sufficient to compute CCE values for evaluation datasets with up to $\approx$ 12,000 examples (which requires, at peak usage, around 10 GB of memory). However, for datasets with significantly larger test splits, modifications to our technique have to be made. One approach to solve this problem is uniform downsampling: select some $k<n$, then from the test set, uniformly sample $k$ input/output pairs. In the notation of Section 3.1.1, if $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}$ is our original sample of test data, then we downsample to obtain $\mathcal{S}_{k}=(x_{j},y_{j})_{j=1}^{k}$.

### v3

@@ A.2 Additional monotonicity experiments
(a) Gaussian Blur ($\tilde{P}_{X}$)
(b) Label noise ($\tilde{P}_{Y}$)
(c) Mixup $\tilde{P}_{Y|X}$.
FG-Net
(d) Gaussian Blur ($\tilde{P}_{X}$)
(e) Label noise ($\tilde{P}_{Y}$)
(f) Mixup $\tilde{P}_{Y|X}$.
Figure 8: Additional perturbation experiments for EVA and FG-Net.
A.3 Using CCE for Machine Learning with Rejection
The rising subfield of “Machine Learning with Rejection” centers around uncovering principled techniques for avoiding predictions when they are expected to be inaccurate or probabilistically misaligned [17]. A typical setup defines some “unreliability score” that we can compute on an arbitrary test input. We then select a threshold, and for any input with an unreliability score above that threshold, we withhold from making a prediction.
A.4 Understanding the behavior of MCMD on known distributions
[…]
Overall, the MCMD appears to behave quite intuitively, with higher values consistently corresponding to areas where the distributions vary from each other more in both mean and standard deviation. Beyond the “eye test”, we specifically highlight a few key desirable characteristics of this metric. First, we see from columns 4 and 5 of Figures 10 and 11 that poor mean fit is penalized more harshly by MCMD when the corresponding variance is low. This indicates that the MCMD requires high variance to explain misalignment of means, which is arguably the correct stipulation to enforce. Additionally, from columns 2 and 3, we observe that the MCMD recognizes the case where the means match but the spread of two distributions is different. This suggests the MCMD is robust at identifying various forms of potential discrepancy between a predictive distribution and the ground truth.
A.5 Additional hyperparameter studies
(a) The effect of $\lambda$ on MCMD computations over synthetic data.
Figure 12: Studying the impact of the regularizer $\lambda$.
Computing the CCE for a given model requires a handful of up-front decisions — primarily, which functions to use for $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ (including kernel hyperparameters) and what values to select for $\lambda$ and $\lambda^{\prime}$. Additionally, one must decide how many Monte Carlo samples, $\ell$, to take when forming $\mathcal{S}^{\prime}$ from a model $f$. In Section 3.2, we provide a variety of “default” settings that have proven effective in our work. To further investigate the impact of the choice of kernel, $\lambda$, and $\ell$, we explore the empirical effects of these hyperparameters. To facilitate our study, we compare samples from two data-generating distributions via MCMD: $(X,Y)$, where $X\sim\texttt{Uniform}(0,2\pi),\text{ }Y|X\sim\mathcal{N}(\cos(X),\frac{1}{4})$ and $(X^{\prime},Y^{\prime})$, where $X^{\prime}\sim\texttt{Uniform}(0,2\pi),\text{ }Y^{\prime}|X^{\prime}\sim\mathcal{N}(X^{\prime}-5,\frac{1}{4})$. We also run a series of CCE computations (with differing hyperparameters) for models trained on COCO-People.
A.5.1 Investigating $\lambda$
To isolate the impact of $\lambda$, we first focus on how it affects the MCMD computation between our samples from $(X,Y)$ and $(X^{\prime},Y^{\prime})$ as specified above. We fix our kernels $k_{\mathcal{X}}$ (RBF kernel with $\gamma=0.5$) and $k_{\mathcal{Y}}$ (RBF kernel with $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$). We vary $\lambda$ along a grid of values between 0 and 1, then compute the MCMD across the input space. For simplicity, we set $\lambda^{\prime}=\lambda$. In Figure 12(a) we plot these MCMD values (second row), along with a scatterplot of the samples being compared (first row). Note that when $\lambda$ is very small, the MCMD becomes quite sensitive to individual points (observe, for example, the spike in MCMD around $2\pi$ for $\lambda=0.001$). As $\lambda$ increases, the MCMD between the two samples grows smoother and tends toward zero. This behavior agrees with Park and Muandet [37]’s categorization of $\lambda$ as a regularizer ($\lambda$ controls the regularization term in the empirical loss Park and Muandet [37] minimize to obtain their estimate of MCMD). Thus, we can consider $\lambda$ as a knob that we can turn depending on how much we “believe” that the samples represent the underlying conditional distributions being compared. When $\lambda$ is large, we discard all evidence from our sample that the distributions are different (thus, the MCMD is 0). When $\lambda$ is small, we place more weight on the individual deviations observed between samples. Generally, $\lambda$ values around 0.05 or 0.1 exhibit reasonable behavior. We set $\lambda=0.1$ in our experiments in the main body of the paper.
We also investigate the effect of $\lambda$ on CCE (which calculates MCMD using samples obtained from a model) for three neural networks trained on the COCO-People dataset. Using the same regressors discussed in Section 4.2, we compute $\overline{\mbox{CCE}}$ across a grid of $\lambda$ values for each model and display the results in Figure 12(b). We observe a similar pattern as before: increasing $\lambda$ shrinks $\overline{\mbox{CCE}}$ toward zero, and we aggressively discard evidence from the data. Interestingly, the relative ordering of the three models is generally preserved across the domain of $\lambda$, suggesting that using $\overline{\mbox{CCE}}$ to make comparisons across models is fairly robust to the choice of $\lambda$.
A.5.2 Modifying $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$
In Section 5, we investigate the impact of the $\gamma$ hyperparameter in $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$, which is present in both the RBF and Laplacian kernels. Here, we provide additional details about this experiment. To obtain our results, we fix $\lambda=0.1$ and vary $\gamma$ simultaneously in $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$. We perform this experiment on the synthetic samples from $(X,Y)$ and $(X^{\prime},Y^{\prime})$ we generated earlier. We use the RBF kernel, but since $\gamma$ induces a similar effect on the Laplacian kernel, we expect an equivalent analysis. Figure 6(a) displays how MCMD computations are affected as $\gamma$ increases. Since the RBF kernel essentially computes similarities by centering a Gaussian at each point with standard deviation $\frac{1}{\sqrt{2\gamma}}$, we translate each $\gamma$ value into an “ellipse of influence” spanning two standard deviations in both the $x$ and $y$ direction. These are visualized in light red around each point in the dataset. As $\gamma$ increases, the number of points that are considered similar to a given $(x,y)$ pair decreases (since the implied Gaussian becomes more concentrated). This increases the sensitivity of the MCMD computation, similarly to selecting low values of $\lambda$. Though most practical applications of MCMD (such as our CCE computation) will not use the RBF kernel for the input values, this behavior provides a good intuition for selecting $\gamma$ for $k_{\mathcal{Y}}$. A workable heuristic is to select $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$, which ensures that the majority of similarities in $\mathcal{Y}$ are nonzero, with closer points exerting more influence.
In the main body, we also compare MCMD values obtained with three different kernels: RBF, Laplacian, and Polynomial (applying the same kernel class to the input and the output spaces). For the RBF and Laplacian kernels, we set $\gamma=2$ for $k_{\mathcal{X}}$ and $\gamma=\frac{1}{2\hat{\sigma^{2}_{Y}}}$ for $k_{\mathcal{Y}}$. For the polynomial kernel, we use a degree of 3, offset of 1, and multiply the inner product by $0.02$ (mostly to ensure MCMD values live on the same scale). For all MCMD computations, we set $\lambda=0.1$. Figure 6(b) displays how the MCMD differs across $\mathcal{X}$ for each kernel. We note that each method largely agrees as to the regions with the most and least discrepancy between the samples. Interestingly, the polynomial kernel produces the smoothest MCMD curve, followed by RBF and Laplacian. From these results, we suspect that the choice of kernel does not matter as much as the hyperparameters associated with that kernel.
A.5.3 Increasing the value of $\ell$
In Section 5, we examine the impact of the number of Monte Carlo samples, $\ell$, on $\overline{\mbox{CCE}}$ (see Figure 7 in the main body). In this experiment, we evaluate the same three models listed in Section 4.4 that were trained on the AAF dataset. For each model we vary the number of samples, $\ell\in\{1,2,3,4,5\}$. We compute $\overline{\mbox{CCE}}$ three times and plot the mean and standard deviation over the three trials. Importantly, we observe that CCE is robust to the number of Monte Carlo samples; this effect is consistent across models. We recommend setting $\ell=1$ to reduce computational complexity, and do so in all of our experiments.
Appendix B Supplementary details
In this section, we provide various details, omitted from the main text for brevity, that we expect will prove useful for reproducibility.
B.1 Derivation of the marginal distribution of $Y$ in Figure 1
In Figure 1, we simulate data from the following data generating process: $X\sim\mathcal{N}(0,1)$, $Y|X\sim\mathcal{N}(\alpha X,1)$, $\alpha=3$, and introduce a poorly specified model, $f$, that ignores the conditional relationship between $X$ and $Y$, but has the same marginal: $f(x)\sim\mathcal{N}(0,1+\alpha^{2})$ .
@@ B.1 Derivation of the marginal distribution of $Y$ in Figure 1
$=\int_{-\infty}^{\infty}\operatorname*{\mathbb{E}}{[Y|X]}P(X=x)dx$ $=\operatorname*{\mathbb{E}}{[\text{Var}(Y|X)]}+\text{Var}(\operatorname*{\mathbb{E}}{[Y|X]})$
$=\int_{-\infty}^{\infty}\alpha xP(X=x)dx$ $=\operatorname*{\mathbb{E}}{[1]}+\text{Var}(\alpha X)$
$=\alpha\int_{-\infty}^{\infty}xP(X=x)dx$ $=1+\alpha^{2}\text{Var}(X)$
@@
B.2 Training specifications
In all experiments where a model was evaluated, we used the training checkpoint associated with the best validation loss for evaluation. This helped mitigate the overfitting that we observed with respect to NLL as training continued. The AdamW optimizer was used for all training jobs, with an initial learning rate of $10^{-3}$ that decayed to 0 on a cosine annealing schedule. We used a weight decay value of $10^{-5}$. For image regression tasks, we applied the standard ImageNet normalization constants and augmented with the AutoAugment [8] transformation during training.
For all datasets, samples were split into train/val/test using a fixed random seed, allocating 70% to train, 10% to val, and 20% to test.
Though the following subsections specify the feature encoder used for each dataset, we refer the reader to our source code  [^2] https://github.com/spencermyoung513/probcal as well as the original literature on each model to find exact details about the regression heads fit on top of each encoder. We now provide further details for the models we trained, organized by dataset. We also list relevant licensing information and URLs for each dataset where applicable.
B.2.1 Synthetic dataset
Following Young et al. [46], the MLPs we trained for the synthetic dataset in Section 4.3 had layers of widths [128, 128, 128, 64]. Training ran for 200 epochs on a 2021 MacBook Pro CPU, with a batch size of 32. We generated this dataset locally using the process defined by Young et al. [46].
B.2.2 FG-Net, AAF, and EVA
@@
B.2.3 COCO-People
On COCO-People, we use the pooled output from a ViT-B backbone (initialized from the vit-base-patch16-224-in21k checkpoint [44, 10]), followed by a [384, 256] MLP with ReLU activations for our feature encoder. This matches the setup used by Young et al. [46]. Models were trained on an internal cluster of 2 NVIDIA GeForce RTX 4090 GPUs with an effective batch size of 256. Training ran for 30 epochs with BFloat16 Mixed Precision. All CCE evaluation was done on a single GPU of this cluster, which has 24GB of VRAM.
The COCO dataset is distributed with a CCBY 4.0 license. See https://cocodataset.org/#home.
Appendix C Computing CCE — practical considerations
In this section, we provide an algorithm for computing CCE, as well as practical considerations that may prove helpful if compute is limited.
C.1 Algorithm to compute CCE for a model
Assuming an existing implementation of the MCMD (which can be easily derived from Equation 2 after a choice of kernel), we employ Algorithm 1 to compute CCE values across a grid $\mathcal{G}\subset\mathcal{X}$ for a given model $f$, embedder $\varphi$, and set of ground-truth samples $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}$.
@@ C.1 Algorithm to compute CCE for a model
$\texttt{S\_prime}\leftarrow\texttt{[]}$ ;
$\texttt{cce\_vals}\leftarrow\texttt{[]}$ ;
@@
for $x$ in $\mathcal{G}$ do
 $\texttt{cce}\leftarrow\texttt{MCMD(S, S\_prime, }\varphi(x)\texttt{)}$ ;
@@
For an implementation of this algorithm in Python, we refer the reader to our source code. The algorithm’s compute is primarily dominated by the matrix multiplications and inversions that are necessary to calculate MCMD, thus time complexity to obtain a single CCE value is $O(n^{3})$, while space complexity is $O(n^{2})$ (assuming, for simplicity, that $n=|\mathcal{S}|=|\mathcal{S^{\prime}}|$).
C.2 Accelerating the CCE computation
Table 3: CCE runtimes for varying values of $n$. For simplicity, it was assumed that $n=|\mathcal{S}|=|\mathcal{S}^{\prime}|=|\mathcal{G}|$. Measurements were taken on a single NVIDIA GeForce RTX 4090 GPU with 24 GB of VRAM.
$n$ 100 1000 3000 5000 10000
Runtime (sec) 0.014 0.224 1.796 5.743 15.406
[…]
The CCE computation can be accelerated even further (and made more numerically stable) by taking advantage of the structure of the Gram matrices involved in the MCMD (see Equation 2). Since $(\mathbf{K}_{X}+n\lambda\mathbf{I}_{n})$ and $(\mathbf{K}_{Y}+m\lambda^{\prime}\mathbf{I}_{m})$ are positive definite and symmetric, we compute their inverses via the Cholesky decomposition (which has half the cost of the typical LU decomposition performed by linear algebra libraries). Our provided code elegantly handles these details, yielding a performant and scalable CCE implementation. Table 3 contains runtimes for our algorithm obtained on a single GeForce RTX 4090 GPU with 24GB of VRAM.
C.3 Addressing memory limitations
As is common with kernel methods, the CCE (which relies on computing the MCMD for a collection of encoded inputs) grows rapidly in terms of memory requirements with the size of the data, exhibiting quadratic spatial complexity. In this work, we found that our hardware was sufficient to compute CCE values for evaluation datasets with up to $\approx$ 12,000 examples (which requires, at peak usage, around 10 GB of memory). However, for datasets with significantly larger test splits, modifications to our technique have to be made. One approach to solve this problem is uniform downsampling: select some $k<n$, then from the test set, uniformly sample $k$ input/output pairs. In the notation of Section 3.1.1, if $\mathcal{S}=(x_{i},y_{i})_{i=1}^{n}$ is our original sample of test data, then we downsample to obtain $\mathcal{S}_{k}=(x_{j},y_{j})_{j=1}^{k}$. It is important to note that none of the experiments from the main body of this paper relied on down-sampling. We present this study solely in the case a practitioner encounters hardware limitations.